OpenStudy (anonymous):

If f(x) = x + 4 and g(x) = 5x - 2, find (g-f)(2).

OpenStudy (priyar):

what is g(x)-f(x) =?

OpenStudy (denonakavro):

Just plug in 2 in for the place of x in both equations f(x) = 2+4 =6 g(x) = 10-2 =8 8-6 = 2 Or you can justify by solving g - f 5x-2-x-4 4x-6 4(2)-6 8-6 = 2

OpenStudy (anonymous):

If f(x) = x + 4 and g(x) = 5x - 2, find (g + f)(-3). could you help with this one?

OpenStudy (denonakavro):

Ok so, Now add: x + 4+5x - 2 6x+2 Now plug in -3 for x -18 + 2 -16 Or we can justify by: Plug in -3 at f(x) -3+4 = 1 Plug in 3 at g(x) -15-2 = -17 -17 + 1 = -16 So the answer is -16

OpenStudy (anonymous):

ok someone showed me the foil way and i got something differ nt

OpenStudy (anonymous):

10. If f(x) = x + 4 and g(x) = 5x - 2, find (gf)(4). and i tried this

OpenStudy (denonakavro):

Yes this is different because this is multiplication, which you will use the FOIL way :)

OpenStudy (denonakavro):

Multiply Using FOIL (x + 4)(5x - 2) 5x^2 -2x +20x -8 Add the similar terms 5x^2 +18x -8 Now plug in 4 as x 5(4)^2 +18(4) -8 80+72-8 182-8 = 174

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

If f(x) = x - 1 and g(x) = 2x - 3, find (f - g)(x). this one i got

OpenStudy (kayders1997):

Did you get an answer?

OpenStudy (kayders1997):

@dulcab526

OpenStudy (anonymous):

yes